Other beamlines: 8.3.1 12.3.1

8.3.1 last updated: Thu Nov 20 19:52:12 PST 2025
sample: no beam mark

camera:

hutch:

no movement for 05:12:18 since Nov 20 14:39:57 2025
beam:


hutch is dark
feedback system is on

last tuneup reached 308% of max at Nov 20 14:25:32 2025
final diode current was 203551 nA, equivalent to 6.18e+11 photons/s at div=2x0.35 mrad, 11110.9 eV and 500 mA (49.07 kGy/s)

cryostream is on at 100K, nozzle in, lN2 level is at 42% and set to autofill
expected to reach 20% at Fri Nov 21 14:58:02 PST 2025

CHL robot dewar is normal, level: 99% and idle, supply idle at 2.69 psi, as of 42 s ago.

current user is AVAILABLE with 12:07:44 left
next user is Ichor Life/Sekar Ramachandran (Josh at Nov 21 08:00:00
door closed for 06:50:29 since Nov 20 13:01:45 2025

00:00:26 ago last control activity at Nov 20 19:51:44 2025
05:13:26 ago at Thu Nov 20 14:38:44 PST 2025
collected: /data/mcfuser/17Nov2025/jamesh/C49/49_1_raster_1_00008.cbf

Exposure_time 0.0990500
Wavelength 0.88557
Detector_distance 300
Start_angle 0.3499
beam stop:

last 10 beam stops:



















/data is 84% full, still room for 1496474 images.

Other beamlines: 8.3.1 12.3.1